s20 608 the end of each year r over its five-year life. In addition to the $20608 cash flow from tions during the fifth and final year, there will be an additional cash flow of $13 200 at the end of the fifth year associated with the salvage value of the system, making the cash fow in yeat 5 cqual to $33 808. Thus, the cash flows associated with this project can be summarised as follows: YEAR CASH FLOW -$54 200 20608 20608 20608 20608 33808 Given a required rate of return of 15%, calculate the following: (a) Payback period (b) Discounted payback period (c) Net present value d) Profitability index (e) Internal rate of return Should this project be accepted?
ST-I In 2010, Caltec Enterprises was considering the acc acquisition of a new conveyor belt st ans. The system required an initial outay f $54200 n Year 0 and had an c or exrpected life of five years The conveyor belt system was expected to reduce the firms material handling costs
Answer #1

Part 1

year

Cash flows

Cumulative cash flows

0

-54200

-54200

1

20608

-33592

2

20608

-12984

3

20608

7624

4

20608

28232

5

33808

62040

Payback period = the year with last negative cumulative cash flows + (7624/20608)

=2+ (7624/20608)

= 2.37 years

Part 2

year

Cash flows

Discounted cash flows (15%)

Cumulative discounted cash flows

0

-54200

-54200

-54200

1

20608

17920 (20608/(1.15^1)

-36280

2

20608

15582.61 (20608/(1.15^2)

-20697.39

3

20608

13550.09 (20608/(1.15^3)

-7147.30

4

20608

11782.69 (20608/(1.15^4)

4635.39

5

33808

16808.55 (33808/(1.15^5)

21443.95

Discounted payback period = the year with last negative Cumulative discounted cash flows + (4635.39/11782.69)

= 3+(4635.39/11782.69)

= 3.39 years

Part 3

year

Cash flows

Present value of cash flows (15%)

0

-54200

-54200

1

20608

17920 (20608/(1.15^1)

2

20608

15582.61 (20608/(1.15^2)

3

20608

13550.09 (20608/(1.15^3)

4

20608

11782.69 (20608/(1.15^4)

5

33808

16808.55 (33808/(1.15^5)

NPV

21443.95

Part 4

Profitability index = 1+(NPV/initial investment) = 1+(21433.95/54200) = 1.40

Part 5

IRR = 29.59%

(Using excel formula IRR)

In excel write all values in consequent rows.

Cash flows

-54200

20608

20608

20608

20608

33808

Then apply the formula,

= IRR (select all values) and then press to get an answer

Yes, the project should be accepted because NPV is positive, profitability index is greater than 1 and IRR is greater than the required rate of return.